# On the dimension distortion under fractionally smooth mappings

### v1

@@ Abstract.
We determine the extent to which continuous mappings in various Sobolev classes distort various dimensions, including the Hausdorff, the upper Minkowski (box-counting), and the upper intermediate dimensions. The intermediate and Minkowski dimension distortion results we obtain are novel already for various classes of fractionally smooth mappings between Euclidean spaces, extending the results of Hencl-Honzík [50, 49] and Huynh [51] to these dimensions. In addition, our work also generalizes the aforementioned results, as well as results of Kaufman [54] and Fraser-Tyson [31], to certain weighted Euclidean spaces and, more generally, to doubling metric measure spaces. As an application of our main result, we quantify a dimension distortion property of quasisymmetric mappings proved by Bishop-Hakobyan-Williams [16] for the intermediate dimension of non-Ahlfors regular subsets of the space.
@@ 1. Introduction
The class of Sobolev mappings has been famously an essential tool in the area of partial differential equations (PDEs) [25]. On the other hand, fractionally smooth Sobolev spaces provide a natural framework for problems where smoothness is intermediate between integer orders, acting as interpolation spaces in the context of functional analysis [1]. They arise naturally in the study of the fractional Laplacian, with applications on anomalous diffusion and jump processes, minimal surfaces, elliptic problems with measure data, and many other areas. We refer to [24] for an even more extensive list of applications and relevant references. In the past two and a half decades, there has been an increasing interest and need to extend this theory of (fractional) Sobolev and, more generally, Triebel–Lizorkin and Besov mappings to metric spaces. Applications of this endeavor include the development of the theory of PDEs [55], calculus of variations [8] and optimal transportation [9] on the non-smooth setting of fractal spaces. The theory of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings defined between metric spaces has been developed by many authors (see, for instance, [17, 38, 40, 41, 56, 62, 64, 57, 42, 63, 34] just to name a few), who have used different approaches to adjust the theory to different settings.
A question of broad interest has been to determine in what ways certain classes of mappings distort dimension notions. One of the earliest results in this direction is by Gehring-Väisälä [33], who gave quantitative bounds on how quasiconformal mappings, a special class of super-critical Sobolev mappings, change the Hausdorff dimension of a subset of $\mathbb{R}^{n}$. Kaufman later proved bounds for the distortion of the Hausdorff and Minkowski dimensions under general super-critical Sobolev mappings [54]. The study of dimension distortion has since been extended to sub-critical Sobolev mappings [48], fractionally smooth Sobolev mappings [50, 49, 51], to other dimension notions [22, 60, 19, 31], and to other settings, such as distortion by Sobolev and quasisymmetric mappings defined between metric spaces [13], [11], [12], [16]. However, in the non-Euclidean setting all results for Sobolev mappings are for the Hausdorff dimension, or cases where all dimensions coincide, with the distortion of the Minkowski dimension only recently settled by the second author [21]. Furthermore, there are currently no results on the dimension distortion under fractional Sobolev, Triebel–Lizorkin, or Besov mappings in the metric setting. In this manuscript, we address both of these open directions by considering the dimension distortion properties of compactly-Hölder mappings, a class that contains and unifies the notions of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings between metric spaces (see Section 2).
@@
Suppose $(X,d_{X})$ is a doubling, uniformly perfect metric space and $(Y,d_{Y})$ is a uniformly perfect metric space. For $p\in(1,\infty)$ and $\alpha\in(0,\infty)$, if $f:X\rightarrow Y$ is $(p,\alpha)$-compactly Hölder and $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)$, then
@@
An immediate corollary is a similar dimension bound under Newtonian and quasisymmetric mappings. In particular, it follows by [21, Theorem 1.2] that a continuous mapping in the Newtonian-Sobolev class is also compactly Hölder for appropriate constants $p$ and $\alpha$, which also yields a similar inclusion for quasisymmetric mappings, under standard assumptions on $X$ and $Y$ (see [21, Corollary 1.3]).
@@
    Suppose $(X,d,\mu)$ is a proper, locally $Q$-homogeneous metric measure space supporting a $Q$-Poincaré inequality for some $Q\in(1,\infty)$, and $(Y,d_{Y})$ is an arbitrary uniformly perfect metric space. Let $f:X\rightarrow Y$ be a continuous mapping with an upper gradient $g\in L^{p}_{\scriptstyle{loc}}(X)$ with $p\in(Q,\infty)$. If $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)<Q$, then
@@
    Suppose $Q\in(1,\infty)$ and $(X,d,\mu)$ is a proper and $Q$-Ahlfors regular metric measure space supporting a $p_{0}$-PI for $p_{0}\in(1,Q)$, and $(Y,d_{Y})$ is a $Q$-Ahlfors regular metric space. Let $f:X\rightarrow Y$ be an $\eta$-quasisymmetric homeomorphism. If $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)\in(0,Q)$, then
@@
Bishop-Hakobyan-Williams [16] studied the quasisymmetric dimension distortion problem problem in the case where the input set $E$ is Ahlfors regular, which implies that all dimensions of $E$ coincide. Their motivation was the absolute continuity on lines property (ACL) that quasisymmetric mappings satisfy in the Euclidean setting. Their result provides a fundamental generalization of this fact in the metric measure spaces setting. In general, however, the Hausdorff, intermediate, and Minkowski dimensions of $E$ could all differ. In such a case, the results from [16] cannot be applied, while (1.3) provides quantitative bounds on $\dim_{\theta}f(E)$, which also recover [21, Theorem 1.2].
@@
Suppose $(X,d_{X},\mu)$ is a locally $Q$-homogeneous metric measure space and $(Y,d_{Y})$ is an arbitrary metric space. Let $s\in(0,\infty)$, $p\in(Q/s,\infty)$, $q\in(0,\infty]$ and $f:X\rightarrow Y$ be a continuous mapping. If $f$ has a finite Hajłasz–Triebel–Lizorkin semi-norm $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$, or a finite Hajłasz–Besov semi-norm $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ with $q\leq p$, then $f$ is (t,sQ/t)(t,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ for $p<q<\infty$, then $f$ is (q,sQ/t)(q,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If, in addition to the assumptions above, $X$ is proper, and $X$ and $Y$ are uniformly perfect, then the following statements hold for any bounded set $E\subset X$ with $\dim_{\theta}E=d_{E}(\theta)<Q$.
@@
    If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ with $p<q<\infty$, then
@@
The above theorem is in fact novel already in the Euclidean case $X=Y=\mathbb{R}^{n}$ with the usual measure and metric, as it is the first intermediate dimension distortion result for classical continuous Triebel–Lizorkin and Besov mappings. In addition, Theorem 1.3 also immediately yields dimension distortion bounds for continuous mappings whose fractional Hajłasz–Sobolev semi-norm $\|\cdot\|_{\dot{M}^{s,p}(X:Y)}$ is finite, since $\dot{M}^{s}_{p,\infty}(X:Y)=\dot{M}^{s,p}(X:Y)$; see Lemma 4.5. Furthermore, if $X,Y$ are not uniformly perfect, similar and simplified arguments that lead to Theorem 1.3 can be employed for $\dim_{0}E=\dim_{H}E$. Consequently, the work of Kaufman on Euclidean super-critical Sobolev mappings [54], the work of Hencl-Honzík on Euclidean Triebel-Lizorkin mappings [50], and the work of Huynh on Euclidean Besov mappings [51] are also recovered. In particular, we have the following unifying result for both the Hausdorff and the Minkowski dimensions.
@@
Although several approaches to defining Sobolev, Triebel–Lizorkin, and Besov spaces on metric spaces have been proposed over the years, a key advantage of the spaces considered in this work is that they support a robust theory without requiring the underlying metric space to be connected. Under additional assumptions on the metric measure space (e.g. Ahlfors regularity, Poincaré inequality), many of the proposed definitions are known to coincide and, hence, Theorems 1.3 and 1.4 can be used to provide dimension distortion bounds for an even broader class of Sobolev, Triebel–Lizorkin, and Besov spaces in that setting. We refer the interested reader to [5, 57, 34] and the references therein for more details.
Note that the bounds involving the Hausdorff and Minkowski dimensions in the above corollary are new for weighted Euclidean spaces, extending the respective results from [50, 49, 51] in that popular setting. For instance, if $\lambda_{n}$ is the $n$-Lebesgue measure, the conditions of Theorem 1.3 (and thus of Theorem 1.4) are satisfied by the weighted Euclidean metric measure space $(\mathbb{R}^{n},d_{\text{euc}},w\lambda_{n})$ for a wide variety of weights $w:\mathbb{R}^{n}\rightarrow[0,\infty]$, such as the class of Muckenhoupt weights (see Chapter 1 in [45]). These weights were introduced by Muckenhoupt [58] in order to characterize the boundedness of the Hardy-Littlewood maximal operator on weighted $L^{p}$ spaces, and have since established an active area within functional and harmonic analysis [36, 37]. More recently, certain Muckenhoupt weights have been associated with dimension theoretic characteristic of spaces through the “weak porosity” notion (see [10, 59]).
This paper is organized as follows. Section 2 includes the necessary background, and introduces an equivalent definition of the intermediate dimension using Hytönen-Kairem dyadic cubes. In Section 3 we prove the intermediate dimension distortion result under compactly Hölder maps (Theorem 1.1). Section 4 contains the proofs of Morrey-type embedding theorems, which we employ to show the Sobolev classes of fractional smoothness in question are contained in the appropriate compactly Hölder class and prove Theorem 1.3. Section 5 contains the proof for the Hausdorff and Minkowski dimension distortion under Hajłasz Triebel–Lizorkin Sobolev and Hajłasz Besov mappings (Theorem 1.4), as well as the proof of Corollary 1.2 for the distortion under Newtonian Sobolev and quasisymmetric mappings.
@@ Acknowledgments.
Part of the project was completed during the first author’s visit to Knoxville and, thus, he wishes to thank the University of Tennessee for their hospitality.
@@ 2. Background
Let $(X,d_{X})$ be a metric space. We often omit the subscript and write $d(x,y)$ for $x,y\in X$ if the space is understood. We denote the open ball centered at $x$ of radius $r>0$ by $B(x,r):=\{z\in X:\,\,d(x,z)<r\}.$ Given a ball $B=B(x,r)\subset X$, we denote by $\lambda B$ the ball $B(x,\lambda r)$, for $\lambda>0$. Given a non-empty set $U\subset X$, we denote by $|U|$ the diameter of $U$ in the metric of $X$. We also make the convention that all bounded sets we consider henceforth are non-empty, even if not explicitly stated, as all results trivially follow otherwise.
@@
This notion is also known as upper box-counting dimension, which justifies the notation with the subscript ‘B’ typically used in the literature (see [26], [29]). We drop the adjective ‘upper’ and the bar notation throughout this paper as we will make no reference to the lower Minkowski dimension. If $X$ is a typical Euclidean space $\mathbb{R}^{n}$, an equivalent formulation for the Minkowski dimension is the following (see [28, Definition 1.1])
@@
Case 2: If $x\in Q_{U}\setminus U$, then $\exists j_{x}\in J^{U}_{k}$ such that
@@
(2.8) $\phi:=\frac{\theta\log\delta_{\varepsilon}}{\log\delta_{\varepsilon}+\log\frac{b\,c_{u}\,c_{0}}{12C_{0}}},$
@@
$\delta_{0}^{\frac{1}{\phi}-\frac{1}{\theta}}=\frac{bc_{u}c_{0}}{12C_{0}},$
which ensures that for all $\delta\leq\delta_{0}$ we have
@@
$\frac{3}{c_{u}c_{0}}\delta^{1/\phi}\leq b^{k_{i}}\leq\frac{\delta}{4C_{0}}$
$b^{k_{i}}\geq$, which makes the collection $\{Q_{i}^{k_{i}}\}$ admissible for $\inf A_{\phi}$.

### v2

@@ Abstract.
We determine the extent to which certain classes of fractionally ‘smooth’ continuous mappings between metric spaces distort various dimensions, including the Hausdorff, upper Minkowski (box-counting), and upper intermediate dimensions. Our intermediate and Minkowski dimension distortion results are new even for continuous (fractional) Sobolev and, more generally, Triebel–Lizorkin and Besov mappings between Euclidean spaces, complementing the work of Hencl-Honzík [52, 51] and Huynh [53]. Moreover, our results also extend the aforementioned work, as well as the work of Kaufman [56] and Fraser-Tyson [33], to certain weighted Euclidean spaces and, more generally, to doubling metric measure spaces. As an application of our main result, we quantify the corresponding dimension distortion properties of quasisymmetric mappings for non-Ahlfors regular subsets of metric measure spaces, strengthening a result of Bishop-Hakobyan-Williams [16].
@@ 1. Introduction
The class of Sobolev mappings has been famously an essential tool in the area of partial differential equations (PDEs) [27]. On the other hand, fractionally smooth Sobolev spaces provide a natural framework for problems where smoothness is intermediate between integer orders, acting as interpolation spaces in the context of functional analysis [1]. They arise naturally in the study of the fractional Laplacian, with applications on anomalous diffusion and jump processes, minimal surfaces, elliptic problems with measure data, and many other areas. We refer to [26] for an even more extensive list of applications and relevant references. In the past two and a half decades, there has been an increasing interest and need to extend this theory of (fractional) Sobolev and, more generally, Triebel–Lizorkin and Besov mappings to metric spaces. Applications of this endeavor include the development of the theory of PDEs [57], calculus of variations [8] and optimal transportation [9] on the non-smooth setting of fractal spaces. The theory of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings defined between metric spaces has been developed by many authors (see, for instance, [17, 40, 42, 43, 58, 64, 66, 59, 44, 65, 36] for a non-exhaustive list), who have used different approaches to adjust the theory to different settings.
A question of broad interest has been to determine in what ways certain classes of mappings distort dimension notions. One of the earliest results in this direction is by Gehring-Väisälä [35], who gave quantitative bounds on how quasiconformal mappings, a special class of super-critical Sobolev mappings, change the Hausdorff dimension of a subset of $\mathbb{R}^{n}$. Kaufman later proved bounds for the distortion of the Hausdorff and Minkowski dimensions under general super-critical Sobolev mappings [56]. The study of dimension distortion has since been extended to sub-critical Sobolev mappings [50], fractionally smooth Sobolev mappings [52, 51, 53], to other dimension notions [23, 62, 19, 33], and to other settings, such as distortion by Sobolev and quasisymmetric mappings defined between metric spaces [13], [11], [12], [16]. Moreover, employing dimension distortion results has yielded applications to the Hölder and quasiconformal classification problems [21, 23, 24], which are often challenging tasks and require the use of heavy machinery. However, in the non-Euclidean setting all results for Sobolev mappings are for the Hausdorff dimension, or cases where all dimensions coincide, with the distortion of the Minkowski dimension only recently settled by the second author [22]. Furthermore, there are currently no results on the dimension distortion under fractional Sobolev, Triebel–Lizorkin, or Besov mappings in the metric setting. In this manuscript, we address both of these open directions by considering the dimension distortion properties of compactly-Hölder mappings, a class that contains and unifies the notions of (fractional) Sobolev, Triebel–Lizorkin, and Besov mappings between metric spaces (see Section 2).
@@
Suppose $(X,d_{X})$ is a doubling, uniformly perfect metric space and $(Y,d_{Y})$ is a uniformly perfect metric space. For $\theta\in(0,1)$, $p\in(0,\infty)$ and $\alpha\in(0,\infty)$, if $f:X\rightarrow Y$ is $(p,\alpha)$-compactly Hölder and $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)$, then
@@
An immediate corollary is a similar dimension bound under Newtonian and quasisymmetric mappings. In particular, it follows by [22, Theorem 1.2] that a continuous mapping in the Newtonian-Sobolev class is also compactly Hölder for appropriate constants $p$ and $\alpha$, which also yields a similar inclusion for quasisymmetric mappings, under standard assumptions on $X$ and $Y$ (see [22, Corollary 1.3]).
@@
    Suppose $(X,d,\mu)$ is a proper, locally $Q$-homogeneous metric measure space supporting a $Q$-Poincaré inequality for some $Q\in(1,\infty)$, and $(Y,d_{Y})$ is an arbitrary uniformly perfect metric space. Let $f:X\rightarrow Y$ be a continuous mapping with an upper gradient $g\in L^{p}_{\scriptstyle{loc}}(X)$ with $p\in(Q,\infty)$. If $\theta\in(0,1)$ and $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)<Q$, then
@@
    Suppose $Q\in(1,\infty)$ and $(X,d,\mu)$ is a proper and $Q$-Ahlfors regular metric measure space supporting a $p_{0}$-PI for $p_{0}\in(1,Q)$, and $(Y,d_{Y})$ is a $Q$-Ahlfors regular metric space. Let $f:X\rightarrow Y$ be an $\eta$-quasisymmetric homeomorphism. If $\theta\in(0,1)$ and $E\subset X$ is bounded with $\dim_{\theta}E=d_{E}(\theta)\in(0,Q)$, then
@@
Bishop-Hakobyan-Williams [16] studied the quasisymmetric dimension distortion problem in the case where the input set $E$ is Ahlfors regular, which implies that all dimensions of $E$ coincide. Their motivation was the absolute continuity on lines property (ACL) that quasisymmetric mappings satisfy in the Euclidean setting. Their result provides a fundamental generalization of this fact in the metric measure spaces setting. In general, however, the Hausdorff, intermediate, and Minkowski dimensions of $E$ could all differ. In such a case, the results from [16] cannot be applied, while (1.3) provides quantitative bounds on $\dim_{\theta}f(E)$, which also recover [22, Theorem 1.2].
@@
Let $(X,d_{X},\mu)$ be a proper locally $Q$-homogeneous metric measure space and $(Y,d_{Y})$ is an arbitrary metric space. Let $s\in(0,\infty)$, $p\in(Q/s,\infty)$, $q\in(0,\infty]$ and $f:X\rightarrow Y$ be a continuous mapping. If $f$ has a finite Hajłasz–Triebel–Lizorkin semi-norm $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$, or a finite Hajłasz–Besov semi-norm $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$ with $q\leq p$, then $f$ is (t,sQ/t)(t,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ for $p<q<\infty$, then $f$ is (q,sQ/t)(q,s-Q/t)-compactly Hölder for all $t\in(Q/s,p)$. If, in addition to the assumptions above, $X$ and $Y$ are uniformly perfect, then the following statements hold for any $\theta\in(0,1)$ and any bounded set $E\subset X$ with $\dim_{\theta}E=d_{E}(\theta)<Q$.
@@
    If $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}<\infty$ with $0<p<q$, then
@@
Theorem 1.3 is in fact novel already in the Euclidean case $X=Y=\mathbb{R}^{n}$ with the usual metric and measure, as it is the first intermediate dimension distortion result for classical continuous Triebel–Lizorkin and Besov mappings, which are known to coincide with the Hajłasz–Triebel–Lizorkin and Hajłasz–Besov mappings in this setting (see Subsection 2.2). In particular, this yields another Euclidean novel dimension distortion result, this time for classical fractional Sobolev mappings. In addition, Theorem 1.3 also immediately yields dimension distortion bounds for continuous mappings whose fractional Hajłasz–Sobolev semi-norm $\|\cdot\|_{\dot{M}^{s,p}(X:Y)}$ is finite, since $\dot{M}^{s}_{p,\infty}(X:Y)=\dot{M}^{s,p}(X:Y)$; see Lemma 4.5. Furthermore, if $X,Y$ are not uniformly perfect, similar and simplified arguments that lead to Theorem 1.3 can be employed for $\dim_{0}E=\dim_{H}E$. Consequently, the work of Kaufman on Euclidean super-critical Sobolev mappings [56], the work of Hencl-Honzík on Euclidean Triebel-Lizorkin mappings [52], and the work of Huynh on Euclidean Besov mappings [53] are also recovered. In particular, we have the following unifying result for both the Hausdorff and the Minkowski dimensions.
@@
Although several approaches to defining Sobolev, Triebel–Lizorkin, and Besov spaces on metric spaces have been proposed over the years, a key advantage of the spaces considered in this work is that they support a robust theory without requiring the underlying metric space to be connected (e.g. Cantor-type sets where other Sobolev notions are trivial), or the measure to be (globally) doubling. Under additional assumptions on the metric measure space (e.g. Ahlfors regularity, Poincaré inequality), many of the proposed definitions are known to coincide and, hence, Theorems 1.3 and 1.4 can be used to provide dimension distortion bounds for an even broader class of Sobolev, Triebel–Lizorkin, and Besov spaces in that setting. We refer the interested reader to [5, 59, 36] and the references therein for more details. It should also be pointed out that we in fact prove the dimension distortion results of Theorems 1.3, 1.4 for mappings with locally finite semi-norms (see Remark 4.9).
Note that the bounds involving the Hausdorff and Minkowski dimensions in Theorem 1.4 are new for weighted Euclidean spaces, extending the respective results from [52, 51, 53] in that popular setting. For instance, if $\lambda_{n}$ is the $n$-Lebesgue measure, the conditions of Theorem 1.3 (and thus of Theorem 1.4) are satisfied by the weighted Euclidean metric measure space $(\mathbb{R}^{n},d_{\text{euc}},w\lambda_{n})$ for a wide variety of weights $w:\mathbb{R}^{n}\rightarrow[0,\infty]$, such as the class of Muckenhoupt weights (see Chapter 1 in [47]). These weights were introduced by Muckenhoupt [60] in order to characterize the boundedness of the Hardy-Littlewood maximal operator on weighted $L^{p}$ spaces, and have since established an active area within functional and harmonic analysis [38, 39]. More recently, certain Muckenhoupt weights have been associated with dimension theoretic characteristics of spaces through the “weak porosity” notion (see [10, 61]).
This paper is organized as follows. Section 2 includes the necessary background for relevant mapping classes and dimensions, and introduces an equivalent definition of the intermediate dimension using Hytönen–Kairema’s dyadic cubes. In Section 3 we prove the intermediate dimension distortion result under compactly Hölder maps (Theorem 1.1). Section 4 contains the proofs of Morrey-type embedding theorems, which we employ to show the Sobolev classes of fractional smoothness in question are contained in the appropriate compactly Hölder class, and prove Theorem 1.3. Section 5 contains the proof for the Hausdorff and Minkowski dimension distortion under Hajłasz Triebel–Lizorkin and Hajłasz Besov mappings (Theorem 1.4), as well as the proof of Corollary 1.2 for the distortion under Newtonian Sobolev and quasisymmetric mappings.
@@ Acknowledgments.
Part of the project was completed during the first author’s visit to Knoxville in Spring 2025 and, thus, he wishes to thank the University of Tennessee for their hospitality.
@@ 2. Background
Let $(X,d_{X})$ be a metric space. We often omit the subscript and write $d(x,y)$ for $x,y\in X$ if the space is understood. We denote the open ball centered at $x$ of radius $r>0$ by $B_{X}(x,r):=\{z\in X:\,\,d(x,z)<r\}.$ If the space is understood, we often omit the subscript $X$. Given a ball $B=B(x,r)\subset X$, we denote by $\lambda B$ the ball $B(x,\lambda r)$, for $\lambda>0$. Given a non-empty set $U\subset X$, we denote by $|U|$ the diameter of $U$ in the metric of $X$. We also make the convention that all bounded sets we consider henceforth are non-empty, even if not explicitly stated, as all results trivially follow otherwise.
@@
This notion is also known as upper box-counting dimension, which justifies the notation with the subscript ‘B’ typically used in the literature (see [28], [31]). We drop the adjective ‘upper’ and the bar notation throughout this paper as we will make no reference to the lower Minkowski dimension. If $X$ is a typical Euclidean space $\mathbb{R}^{n}$, an equivalent formulation for the Minkowski dimension is the following (see [30, Definition 1.1])
@@
Case 2: If $x\in Q_{U}\setminus U$, then there is $j_{x}\in J^{U}_{k}$ such that
@@
(2.8) $\phi:=\frac{\theta\log\delta_{\varepsilon}}{\log\delta_{\varepsilon}+\theta\log\frac{b\,c_{u}\,c_{0}}{12C_{0}}},$
@@
$\delta_{\varepsilon}^{\frac{1}{\phi}-\frac{1}{\theta}}=\frac{bc_{u}c_{0}}{12C_{0}},$
which ensures that for all $\delta\leq\delta_{\varepsilon}$ we have
@@
$\frac{3}{c_{u}c_{0}}\delta^{1/\phi}\leq b^{k_{i}}\leq\frac{\delta}{4C_{0}},$
which makes the collection $\{Q_{i}^{k_{i}}\}$ admissible for $\inf A_{\phi}$.